40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Dec 9 19:19:04 2019, Yehonathan, Update, PSL, AOM first order beam alignment 20191209_193112.jpg20191209_193203_HDR.jpgimageedit_2_7551928142.gifimageedit_3_5863650538.gif
    Reply  Tue Dec 10 01:24:17 2019, Yehonathan, Update, PSL, AOM first order beam alignment 
       Reply  Tue Dec 10 13:26:46 2019, Yehonathan, Update, PSL, AOM first order beam alignment 
    Reply  Fri Jan 17 21:56:22 2020, Yehonathan, Update, PSL, AOM first order beam alignment 20200117_174841.jpg
       Reply  Wed Jan 22 11:00:21 2020, gautam, Update, PSL, PMC REFL ghost beam 
Message ID: 15087     Entry time: Mon Dec 9 19:19:04 2019     Reply to this: 15089   15131
Author: Yehonathan 
Type: Update 
Category: PSL 
Subject: AOM first order beam alignment 

{Yehonathan, Rana}

In order to setup a ringdown measurement with perfect extinction we need to align the first order beam from the AOM to the PMC instead of the zeroth order.

We connected a signal generator to the AOM driver and applied some offset voltage. We spot the first order mode and align it to the PMC. The achieved transmitted power is roughly as it was before this procedure.

Along the way few changes has been made in the PSL table:

1. Some dangling BNCs were removed.

2. Laser on the south east side of the PSL table was turned off.

3. DC power supplies were removed (Attachment 1 & 2). The rubber legs on the first one are sticky and leave black residue.

4. The beam block that orginally blocked the AOM high order modes was raised to block the zeroth order mode (Attachment 3).

5. The unterminated BNC T junction (Attachment 4 - before picture). from the PMC mixer to the PMC servo was removed.

However, we are currently unable to lock the PMC on high gain. When the gain is too high the PZT voltage goes straight to max and the lock is lost.

Attachment 1: 20191209_193112.jpg  225 kB  Uploaded Mon Dec 9 20:20:08 2019  | Hide | Hide all
20191209_193112.jpg
Attachment 2: 20191209_193203_HDR.jpg  282 kB  Uploaded Mon Dec 9 20:20:20 2019  | Hide | Hide all
20191209_193203_HDR.jpg
Attachment 3: imageedit_2_7551928142.gif  989 kB  Uploaded Mon Dec 9 20:20:40 2019  | Hide | Hide all
imageedit_2_7551928142.gif
Attachment 4: imageedit_3_5863650538.gif  1.368 MB  Uploaded Mon Dec 9 20:22:23 2019  | Hide | Hide all
imageedit_3_5863650538.gif
ELOG V3.1.3-